# Ayoung Lee1, Ryan Sungmo Kwon1, Peter Railton2, Lu Wang1

### v2

@@ 4 Experiments and Analyses
In this section, we first present the main results for decision ambivalence, psychological discomfort, and value shifts in perspectives (Section 4.1). We then discuss the results of the steerability analysis using different value frameworks (Section 4.2). Finally, we examine the impact of question framing from first-person perspectives vs. third-person perspectives in Section 4.3. The prompts we use are presented in Appendix A.2.4. We generate both answers and rationales for the questions. In this section, we use the answers, while the rationales are used for the error analysis (Appendix A.4.2).
@@ 4.1 Main Results
We evaluate both open and closed models, including Qwen-2.5 7B and 72B (Yang et al., 2024), Llama-3.1 8B (Dubey et al., 2024) and 3.3 70B (AI, 2024a), Mistral 24B (AI, 2024c) and 123B (AI, 2024b), GPT-4o and its mini version (Hurst et al., 2024), Claude Sonnet v2 and Haiku v1 (Anthropic, 2024). We set the temperature to zero to ensure greedy decoding. The results are summarized in Table 1. Overall performance averaged over all categories indicates that Claude Sonnet achieves the highest accuracy at 76.76%.
We also focus on three key dimensions: Ambivalence, Discomfort, and Value Shifts, as defined in Section 1. For Ambivalence, we use the general question in Simple Contrast and Swayed Contrast. For Discomfort, we use the discomfort question in Straightforward and Swayed Contrast. For Value Shifts, we use the previous and current questions in Half-Shift and False-Shift. Each of the categories include two character descriptions except for Simple Contrast, so we average results across both descriptions when applicable. The findings show that large proprietary models, such as Claude Sonnet and GPT-4o, generally outperform others across most settings. However, the open-weight Mistral models (Mistral-24B and Mistral-123B) uniquely outperform all other models in settings where the ground-truth answer is ”Ambiguous” (Simple Contrast, Half-Shift), pointing to limitations in closed models’ performance on ambivalent situations. A Mann-Whitney U test (McKnight & Najab, 2010) confirmed the statistical significance of this observation ($p<0.05$). The complete results across all questions and settings are shown in Figure A.4.1.
@@
Model Overall Prev. Curr. Prev. Curr.
Qwen2.5-7B 59.61 35.65 70.00 61.45 81.59 64.35 37.97 69.57 57.10
Qwen2.5-72B 74.07 38.26 92.03 76.38 81.34 96.23 47.97 89.71 72.90
Llama3.1-8B 59.97 43.19 60.58 52.46 87.06 69.60 42.73 70.88 45.86
Llama3.3-70B 75.63 54.78 84.20 91.16 87.25 96.81 43.19 89.13 66.52
Mistral-24B 74.24 61.74 78.70 91.16 57.87 96.67 61.59 92.46 64.64
Mistral-123B 76.45 63.19 84.78 80.72 87.83 96.81 65.07 90.87 59.57
GPT-4o-mini 67.37 26.45 86.34 76.74 78.63 90.99 24.56 83.43 50.73
GPT-4o 75.87 38.37 92.73 96.51 71.95 98.84 55.52 93.46 66.72
Claude-Haiku 70.71 53.33 87.10 77.97 65.37 96.96 56.03 89.13 31.88
Claude-Sonnet 76.76 49.28 93.91 82.46 88.41 99.13 43.62 91.16 76.52
@@
Table 1: Results of accuracy for understanding ambivalence, discomfort, and value shift for closed (upper) and open-weight (lower) models. Greedy decoding is used, and the best performance per column is indicated as bold. For value shift, accuracy on previous vs. current questions is compared, with higher values highlighted in blue. The random baseline is 0.50 for the discomfort question and 0.33 for the other questions. Models perform poorly on predicting ambivalence and show notable drops on current questions, indicating difficulty with dynamic value shifts.
@@
RQ1: Do LLMs understand ambivalence involved in the decision-making process? We test whether models incorrectly respond with ”Ambiguous” when a clear ”Yes” or ”No” is needed, and vice versa. This distinction matters: in clear-cut scenarios, ”Ambiguous” is unhelpful, while in truly ambivalent cases, a definitive answer can be misleading. We present two takeaways based on the results.
1) Models struggle to understand ambivalence. As demonstrated by the relatively low accuracy in the Simple Contrast category, even advanced models like GPT-4o and Claude-Sonnet failed to surpass 60%, with the highest performance being 63.19% by Mistral-123B. Notably, models in the Mistral family consistently outperformed both open-weight and closed alternatives in Simple Contrast category, challenging conventional assumptions about the superiority of proprietary models in value reasoning. 2) Performance diverges across settings, revealing a trade-off between clear-cut and ambivalent decisions. GPT-4o and Claude-Sonnet, both closed models, achieved the top two scores in the Swayed Contrast setting, while open-source models from the Mistral family outperformed others in the Simple Contrast setting. This divergence reveals that no single model consistently excels across both the Simple Contrast and Swayed Contrast categories, suggesting a trade-off: models proficient in identifying ambivalent situations are more likely to output ”Ambiguous” even in cases that are clear-cut, and conversely, models that perform well in clear-cut settings exhibit reduced accuracy in recognizing ambivalence. Future research could focus on aligning models’ standards for identifying ambivalence with human judgments, as current models are either too strict or too lenient compared to human expectations.
@@
RQ2: Do LLMs exhibit discomfort when making hard decisions? Recall that in the Straightforward category, the character firmly supports one value over the other, so the answer to the discomfort question should be ”No.” Conversely, in the Swayed Contrast category, the character holds a priority but acknowledges both values, making ”Yes” the appropriate response. Most models exhibited strong overall accuracy in predicting discomfort, with GPT-4o achieving the highest accuracy of 96.51% in the Straightforward category, and Claude-Sonnet reaching 88.41% in the Swayed Contrast category. Moreover, performance declined when predicting responses to the discomfort question in the Straightforward category when the ground-truth answer was ”No” compared to when it was ”Yes.” We hypothesize that Straightforward characters have extreme perspectives, leading them to prefer taking action over doing nothing. Refer to Appendix A.4.3 for further results and discussion.
@@
RQ3: Can LLMs reason about value shifts? We challenge the LLMs to detect the existence of value change within the perspectives. Recall that dynamic characters initially follow a single value-based rationale but experience a transformation in their value systems (Figure 2). Therefore, answering the previous question requires simpler reasoning than answering the current question correctly. This difference in complexity is reflected in model performance. Accuracies for the previous questions are generally high across models, with all models except for smaller ones under 10B parameters (Qwen2.5-7B, Llama3.1-8B) achieving above 89%. However, as expected, there is a consistent performance drop on the current question across all models. The Wilcoxon test (Woolson, 2005) proves the statistical significance of the drop ($p<0.0001$). These results suggest that models struggle to adapt to the updated values, possibly because they fail to detect the value change, do not fully understand the more complex updated values, or both.
4.2 Steerability Analysis Using Diverse Value Frameworks
@@
Our value-related rationales are specific to the situation. To facilitate a more generalized analysis, in this section, we map these rationales onto broader value frameworks and subsequently conduct an analysis on LLM steerability based on value pairs. Following DailyDilemmas (Chiu et al., 2024), we use four value frameworks: the World Values Survey (WVS, 2023-02-17), Moral Foundations Theory (Graham et al., 2013), Maslow’s Hierarchy of Needs (Maslow, 1943), and Aristotle’s Virtues (Grant et al., 1874). In CLASH, each dilemma includes two competing value-related rationales, which we map to the 301 intermediate values defined in DailyDilemmas, such as Justice, Autonomy, or Loyalty. DailyDilemmas also provides predefined mappings from these intermediate values to dimensions in the four value frameworks. For example, Justice is mapped to Fairness in Moral Foundations Theory and Safety in Maslow’s Hierarchy of Needs. By applying these mappings, we obtain the value dimensions associated with each side of the dilemma and use them for our analysis. Further details regarding the mapping process can be found in Appendix A.4.4.
Dataset Distribution
Figure 3: Distribution of value dimensions. Percentages represent the proportion of dilemmas in which each dimension appears in at least one of the two competing rationales.
Figure 3 shows the percentage of each value dimension in our dataset. Based on the acquired mappings, we calculate the coverage of each value dimension, defined as the percentage of dilemmas in which the dimension appears in at least one of the two competing rationales. Our dataset demonstrates relatively high and balanced coverage across all dimensions of the World Values Survey, with each dimension exceeding 75%. In terms of the Moral Foundations Theory, the dimensions of Fairness, Authority, and Care exhibit substantially higher coverage (each above 75%) compared to Loyalty and Purity, which show notably lower coverage (each below 45%). Regarding Maslow’s Hierarchy of Needs, the upper four levels (Self-Esteem, Safety, Self-Actualization, and Love and Belonging) are well represented, each with coverage exceeding 75%. In contrast, the lowest level, Physiological, is less frequently represented in real-life dilemmas, with a coverage rate of only 10.23%. For Aristotle’s Virtues, the distribution is notably uneven. Truthfulness exhibits the highest coverage at 60.82%, whereas Modesty and Patience are scarcely represented, with coverage rates of just 0.58% and 0.29%, respectively.
Measuring Steerability
Model Correlation P-value
Qwen -0.634 2.90e-4
Llama -0.675 8.25e-5
Mistral 0.136 4.90e-1
GPT-4o 0.156 4.27e-1
Claude -0.322 9.47e-2
Overall -0.329 7.27e-5
Table 2: Spearman’s correlation coefficient between preference and steerability. A significant negative correlation was observed for Qwen, Llama, and all models combined.
We first identified pairs of dimensions from each framework. For example, Moral Foundation Theory encompasses five dimensions: Care, Fairness, Purity, Loyalty, and Authority. This results in ten combinations of competing value dimension pairs, derived from selecting two out of five options. An exception to this general approach is the World Values Survey, which defines only two value pairs: Survival vs. Self-Expression and Traditional vs. Secular-Rational. For each competing pair, we initially filter all situations explicitly associated with the respective pair. For example, in examining the competing value pair Safety vs. Self-Esteem, we specifically select situations in which Safety is related to the acceptable rationale and Self-Esteem is related to the unacceptable rationale, or vice versa. We then calculate three preferences: (i) the base preference by prompting without any character description, (ii) preference after being steered towards Safety using one of the Swayed Contrast character description, and (iii) the preference after being steered towards Self-Esteem using the other Swayed Contrast character description. The specific methodology for calculating preferences is detailed in Appendix A.4.5, Algorithm 1. We assign numbers to the responses ”Yes,” ”No,” and ”Ambiguous,” sum these numbers, and then compute their average to obtain each preference. Then we use the preferences to measure steerability, with the detailed procedure elaborated in Appendix A.4.5.
Findings
@@
We first calculated the Spearman’s correlation coefficient (Spearman, 1961) between preference and steerability to understand whether inherent tendencies affect flexibility and provide insights into strategies for increasing steerability. We hypothesize that strong inherent preference for one value over the other within a pair would hinder the ability to steer towards the opposing value. We collect base preference and difference of steerability across all value pairs and models and calculated the correlation (Table 2). Our analysis revealed a significant negative correlation ($r=-0.329,p<0.0001$), confirming our hypothesis and suggesting that greater preference indeed restricts steerability. When considering individual model families, we observed nuanced results, which is in Table 2. Specifically, Qwen and Llama demonstrated a statistically significant negative correlation ($p<0.001$), while Mistral, GPT-4o, and Claude did not exhibit statistically significant correlations. We recommend further data collection and comprehensive analysis to elucidate the correlation dynamics for the Mistral, GPT-4o, and Claude model families.
We also investigate how steerability varies between smaller and larger models. To do this, we conduct Wilcoxon test (Woolson, 2005) comparing small and large models, focusing on two aspects: (1) the absolute difference in steerability between the two values within each pair, and (2) the aggregate steerability, which is the sum of steerability towards each value in a pair. For example, consider the value pair ”safety vs. self-esteem.” If steerability toward ”safety” is 0.7 and steerability toward ”self-esteem” is 0.8, the absolute difference in steerability is 0.1, and the total steerability is 1.5.
@@
The first Wilcoxon test revealed that smaller models (Qwen2.5-7B, Llama3.1-8B, Mistral-24B, GPT-4o-mini, Claude-Haiku) are significantly more likely to show greater steerability toward one value over the other ($p<0.001$), whereas the second Wilcoxon test revealed that larger models (Qwen2.5-72B, Llama3.3-70B, Mistral-123B, GPT-4o, Claude-Sonnet) exhibit significantly higher overall steerability ($p<0.0001$). Per-value steerability comparison results can be found in Appendix A.4.6.
4.3 Question Framing Comparisons via First-Person vs. Third-Person Perspectives
Model Straight Simple Swayed
Qwen2.5-7B -1.74 8.12 -0.58
Qwen2.5-72B 0.00 -1.45 -2.46
Llama3.1-8B 5.80 -6.67 4.78
Llama3.3-70B -0.58 -7.83 -0.87
Mistral-24B -1.01 0.00 -7.25
Mistral-123B -2.17 -6.09 -4.49
GPT-4o-mini -0.73 -6.10 -1.02
GPT-4o -0.15 -2.33 -3.20
Claude-Haiku -0.72 -0.29 -8.84
Claude-Sonnet -0.14 -14.78 -1.16
Table 3: Accuracy difference between first-person and third-person. Red cells are below zero, indicating worse performance by the 1st party setting. Blue cells indicate better performance in the 1st party setting.
We conduct a comparative analysis of question framing to assess the influence of first-person versus third-person perspectives on the steerability of LLMs. Drawing on the concept of ”self-distancing” (Grossmann & Kross, 2014), which shows that people reason more wisely from a third-person perspective, we hypothesize that first-person framing may make scenarios feel more personally relevant to the model, reducing its objectivity. Thus, we expect LLMs to be more resistant to value steering when questions are framed in the first-person perspective. Refer to Appendix A.2.4 for the specific prompts. Indeed our findings indicate a notable decline in performance across nearly all experimental conditions except for Llama3.1-8B (Table 3) when first-person perspectives are present. This performance degradation is confirmed as statistically significant by a Wilcoxon test (Woolson, 2005) ($p<0.0001$). The drop in accuracy was more pronounced in the Simple Contrast and Swayed Contrast categories compared to the Straightforward category. Notably, Claude-Sonnet exhibited the largest performance decrease of 14.78% in Simple Contrast, while Claude-Haiku showed the greatest drop of 8.84% in Swayed Contrast.
Next, we examine differences in steerability according to specific values and their opposing counterparts. The analysis revealed that most value pairs exhibited increased steerability when framed from a third-person perspective. However, certain cases colored green in Figure 4 demonstrated greater effectiveness when employing a first-person perspective. By identifying the most effective framing perspective for each value pair, we aim to present patterns that may inform strategies for enhancing steerability in different value scenarios. The results of this analysis are illustrated in Figure 4. Specifically, the first-person perspective significantly improved steerability for the following value scenarios: Safety when Self-Esteem is presented as the competing dimension, Fairness when Purity is presented as the competing dimension, and Fairness when Care is presented as the competing dimension. The setting of Truthfulness when Friendliness is presented as the competing dimension showed negligible differences between perspectives. For all other examined value pairs, the third-person perspective consistently demonstrated superior effectiveness in enhancing steerability.
[…]

### v3

@@ 4 Experiments and Analyses
We first present the main results for ambivalence, discomfort, and value shifts (§4.1), followed by an in-depth analysis of the reasoning chains thinking models produce (§4.2). Then, we analyze conditional steerability of models across value frameworks (§4.3). Finally, we compare prompting strategies, contrasting first- vs. third-person question framing (§4.4). For all experiments, we prompt the LLMs to produce their reasoning process followed by the answer, with the full prompt given in Appendix C.2.
@@ 4.1 Main Results
We evaluate both general-purpose LLMs and reasoning language models (RLMs; also called thinking models). Specifically, we include five families of LLMs, each with two sizes: Qwen-2.5 (Yang et al., 2024), Llama-3 (Dubey et al., 2024; AI, 2024a), Mistral (AI, 2024c; b), GPT-4o (Hurst et al., 2024), Claude-3.5 (Anthropic, 2024), and four RLMs: Qwen3-32B (Yang et al., 2025), GPT-5 (OpenAI, 2025), Claude-4 Sonnet (Anthropic, 2025), Deepseek-3.1 (Liu et al., 2024). Refer to Appendix C.1 for the detailed list of models. We set the temperature to 0.0 for greedy decoding.
The main results are summarized in Table 1, with more detailed results broken down into questions and settings in Table 5 and Table 6222We recruited two human annotators who had no knowledge of the research context to provide responses to 50 randomly sampled situations. The overall accuracy of humans was 92.8. Discussions with them confirmed that errors arise from the inherent complexity of the task rather than the limitations of the design.. Claude-4 Sonnet, a RLM, achieved the highest accuracy across all categories (88.89). Overall, within each family (Qwen, GPT, and Claude), RLMs consistently outperformed their LLM counterparts, indicating that enhanced reasoning capabilities benefits value understanding and dilemma interpretation. In particular, the average output length is 674.5 tokens for RLMs, compared to that of 142.8 for LLMs. We also provide an examination of reasoning characteristics (§4.2) and an error analysis (Appendix D.2) using the reasons produced. The remainder of this section focuses on three key dimensions: Ambivalence, Discomfort, and Value Shifts.
@@
Model Overall Prev. Curr. $\Delta$ Prev. Curr. $\Delta$
Non-thinking models
Qwen2.5-7B 65.65 42.61 68.84 59.95 80.77 59.13 43.49 15.64 67.83 53.14 14.69
Qwen2.5-72B 82.75 38.26 91.74 79.33 80.58 95.65 48.99 46.66 89.28 71.01 18.27
Llama3.1-8B 65.90 51.59 54.06 56.43 80.30 72.46 46.31 26.15 70.43 42.03 28.40
Llama3.3-70B 83.28 51.88 82.75 91.15 85.37 96.23 42.75 53.48 90.00 65.65 24.35
Mistral-24B 81.19 62.03 81.01 88.25 55.14 96.81 61.16 35.65 91.30 66.96 24.34
Mistral-123B 85.09 62.90 85.22 86.61 86.21 96.81 68.41 28.40 89.42 60.58 28.84
GPT-4o-mini 75.47 26.45 86.34 76.80 75.83 90.99 24.56 66.43 83.43 50.73 32.70
GPT-4o 84.75 38.37 92.73 96.50 70.59 98.84 55.52 43.32 93.46 66.72 26.74
Claude-3.5-Haiku 77.86 53.33 87.10 78.48 60.60 96.96 56.03 40.93 89.13 31.88 57.25
Claude-3.5-Sonnet 85.37 49.28 93.91 83.16 87.62 99.13 43.62 55.51 91.16 76.52 14.64
Thinking models
Qwen3-32B 84.40 36.84 94.45 94.69 91.02 90.00 42.32 47.68 91.01 77.34 13.67
GPT-5 86.14 24.06 95.36 96.35 98.61 97.54 59.86 37.68 81.45 77.39 4.06
Claude-4-Sonnet 88.89 51.01 95.22 92.69 94.58 98.55 51.30 47.25 92.17 86.96 5.21
Deepseek-3.1 84.54 12.46 96.96 94.23 94.70 93.91 37.39 56.52 90.72 87.10 3.62
@@
Table 1: Accuracy for understanding ambivalence, discomfort, and value shift. Results are averaged over two character descriptions per category except for Simple Contrast that contains only one description. Greedy decoding is used. The best performance per task is bolded, with the second-best underlined. For value shift, darker blue shade indicates bigger accuracy difference between previous and current questions. A random baseline is 0.50 for the discomfort question and 0.33 for the other questions. Models perform poorly on ambivalence prediction and show notable drops on current questions, indicating difficulty with interpreting dynamic value shifts.
@@
RQ1: Do LLMs understand the ambivalence involved in the decision-making process? We employ the general question in Simple Contrast and Swayed Contrast to test whether models erroneously return “Ambiguous” when a clear “Yes” or “No” is needed, and vice versa. This distinction matters: in clear-cut scenarios, “Ambiguous” is uninformative, while in truly ambivalent cases, a definitive answer can be misleading and even lead to unrecoverable consequences in high-stakes contexts. We present three key takeaways as follows.
1) Models struggle to understand ambivalence. As shown by the relatively low accuracy in the Simple Contrast category, models struggle to recognize ambivalence, with Mistral-123B achieving the highest accuracy of only 62.90. 2) Models trade off between clear-cut and ambivalent decisions, potentially exposing limitations in model training. GPT-5 and Deepseek-3.1 achieve the highest scores in the Swayed Contrast setting, but open-weight models from the Mistral family outperform others in Simple Contrast. No model excels in both, challenging conventional assumptions about the superiority of proprietary models in value reasoning (Nie et al., 2023; Feng et al., 2024). 3) RLMs outperform in clear-cut cases while underperform in ambivalent settings. Within the same family, models that use explicit thinking modes generally perform worse on ambivalent cases but better on clear-cut ones. An exception is Claude-4 Sonnet, for which the accuracy increased in all ambivalent cases. To gain insights, we compare the reasoning chains of Claude-4 Sonnet with other models’ thoughts. For Qwen3-32B and Deepseek-3.1, the reasoning chains often include hedging expressions such as “might be acceptable”, yet still produce definitive answers, indicating limited recognition of ambivalence (Figure  3). In contrast, the reasoning of Claude-4 Sonnet is closely aligned with its answers: uncertain wording with ambivalent answers or decisive wording with clear-cut ones. More examples of this phenomenon can be found in Appendix E.1.
Reasoning: Character A would have to balance these two values. If they debit the account, they’re violating the customer’s consent, which is a big deal. But if they don’t, Carole loses her job. Since the customer didn’t consent, and the action is against their principles, maybe the answer is no. But maybe they see it as a necessary evil.
Answer: No
Figure 3: Sample reasoning by Qwen3-32B. Hedging expressions do not align with the incorrect, definitive answer of “No”.
@@
RQ2: Do LLMs exhibit discomfort when making hard decisions? Recall that in the Straightforward category, characters strongly favor one value, so the discomfort answer is “No.” In Swayed Contrast, characters hold a priority but acknowledge both values, making “Yes” the appropriate response. We evaluate the discomfort question for both categories. Most models exhibit strong abilities in discomfort understanding, with GPT-4o achieving the highest accuracy of 96.50 in the Straightforward category, and GPT-5 reaching 98.61 in the Swayed Contrast category. Further analysis of differences in discomfort questions (Appendix D.3) reveal that Straightforward characters, due to their extreme perspectives, prefer action over inaction.
@@
RQ3: Can LLMs reason about value shifts? We challenge LLMs to detect the existence of value change within the perspectives, using Half-Shift and False-Shift categories. Recall that dynamic characters initially follow a single value-related rationale but experience a transformation in their value systems (Figure 2). Therefore, answering the previous question requires simpler reasoning than answering the current question. As anticipated, we observe a consistent performance drop in the current question across all models (Wilcoxon test (Woolson, 2005), $p<0.0001$), with the largest accuracy drop of 66.43 points for GPT-4o-mini and an average decrease of 42.95 points. These results suggest that models struggle to adapt to the updated values, possibly due to the incapability of detecting value change and grasping the associated complexity.
4.2 Further Reasoning Chain Analyses
Building on the success of RLMs in §4.1, we further analyze the reasoning chains RLMs produce to identify successful characteristics and provide insights for advancing value reasoning in LMs.
(a) Cognitive characteristics of reasoning. Successful chains use less backtracking and verification, suggesting sometimes they are unhelpful for value reasoning. The failure chains contain more early- and overcommitment.
(b) Usage of ethical theories in reasoning. Successful chains include more of pragmatic ethics and rights ethics, showing the importance of real-life adaptation.
Figure 4: Characteristics of successful vs. unsuccessful reasoning chains.
Cognitive Characteristics.
Inspired by the math reasoning analysis done in Gandhi et al. (2025), we start with probing four cognitive characteristics of the thoughts: backward chaining, verification, backtracking, and subgoal setting, with the goal of evaluating whether any of these aid value reasoning. Initial exploration suggests that backward chaining and verification are more prevalent and relevant to the problems in CLASH, which will be the focus of this study. The results, shown in the first column of Figure 4(a) (prompts in Appendix E.2), reveal that in contrast to math and game tasks, successful value-laden reasoning exhibits less backward chaining and verification.
After taking a closer inspection of unsuccessful reasoning chains, we find two patterns: early commitment and overcommitment of incorrect judgments, and these are especially pronounced in ambivalence and discomfort prediction. When judging decision ambivalence, some erroneous chains prematurely favor one side (early commitment) and then persistently pursue that side (overcommitment). When it comes to discomfort prediction, certain chains hastily conclude that the character would feel discomfort (early commitment) and redirect to the general question, focusing on whether the action is acceptable (overcommitment). Illustrative examples are provided in Appendix E.3, with associated prompts in Appendix E.2. A systematic analysis shown in the second and third columns of Figure 4(a) reveals that early commitment and overcommitment occur more frequently in unsuccessful chains.
Usage of Ethical Theories.
Ethical theories provide a normative structure for reasoning, distinct from value frameworks. Using the theories outlined in Chakraborty et al. (2025), i.e., Care Ethics (Gilligan, 1993), Deontology (Alexander & Moore, 2007), Ethical Pluralism (Ross, 2002), Pragmatic Ethics (Dewey & Tufts, 2022), Rights Ethics (Dworkin, 2013), Utilitarianism (Mill, 2016), and Virtue Ethics (Hume, 2000), we analyze successful and unsuccessful reasoning chains based on how frequently these theories are applied. Figure 4(b) shows that models most often rely on deontology, utilitarianism, and virtue ethics. When compared with unsuccessful chains, successful chains demonstrate greater emphasis on pragmatic and rights-based ethics, with pragmatic ethics especially prominent given its focus on real-world adaptation. This points to future work that explores strategies to better elicit successful chains grounded in ethical theories.
4.3 Conditional Steerability Analysis Using Diverse Value Frameworks
@@
Our value-related rationales are specific to the situation. To enable a generalized analysis, we map the rationales to broader value frameworks and subsequently conduct an analysis on LLM steerability. In particular, we propose a method for evaluating conditional steerability: steerability towards one value over the other. This setting considers more complexities encountered in real life than prior work (Dong et al., 2023; Rimsky et al., 2024), which has mainly focused on alignment with a single dimension without considering its opposing counterpart. Following DailyDilemmas (Chiu et al., 2024), we use four value frameworks: the World Values Survey (WVS, 2023-02-17), Moral Foundations Theory (Graham et al., 2013), Maslow’s Hierarchy of Needs (Maslow, 1943), and Aristotle’s Virtues (Grant et al., 1874). Remember that each dilemma in CLASH includes two competing value-related rationales; we map each rationale to the 301 intermediate values defined in DailyDilemmas (e.g., Justice, Autonomy), then map these values further onto dimensions from four broader value frameworks. [^3] DailyDilemmas also provides predefined mappings from the 301 intermediate values to dimensions in the four value frameworks. For example, Justice is mapped to Fairness in Moral Foundations Theory. Using the mappings, we identify the competing value dimension pairs. Each side may be linked with multiple value dimensions, and each dilemma may involve multiple competing value pairs. More details on the mapping process can be found in Appendix F.3, while Appendix A.4 presents the dataset statistics of the mapped values.
Measurement of Conditional Steerability.
We first identify pairs of dimensions for each framework. For instance, Moral Foundation Theory encompasses five dimensions: Care, Fairness, Purity, Loyalty, and Authority. This results in ten combinations of competing value dimension pairs. [^4] An exception to this general approach is the World Values Survey, which defines only two value pairs (WVS, 2023-02-17): Survival vs. Self-Expression and Traditional vs. Secular-Rational.
Figure 5: Illustration of conditional steerability measurement. A model perfectly steered toward Safety is assigned a value of 0, while one perfectly steered toward Self-Esteem is assigned a value of 1. The three preferences are obtained by prompting LLMs and collecting responses of “Yes”, “No”, and “Ambiguous”. a, b, c, and d are computed as the differences between the corresponding elements shown in the image.
For each competing pair, we initially filter all situations associated with the respective pair. For example, in examining the competing value pair Safety vs. Self-Esteem, we select situations in which Safety is mapped to the supporting rationale and Self-Esteem is mapped to the opposing rationale or vice versa. [^5] To ensure reliable results, we only consider value pairs that occur more than 16 times, representing around the top 25% of all pairs. We then calculate three preferences: (i) the base preference by prompting without any character description, (ii) preference after being steered towards Safety using one of the Swayed Contrast character description, and (iii) the preference after being steered towards Self-Esteem using the other Swayed Contrast character description. We assign numerical values to the responses to the general question (“Yes”: 1 or 0, “No”: 0 or 1 and “Ambiguous”: 0.5). Then, the occurrence of each response type is tallied and normalized by the total number of questions to obtain the preference score. The specific methodology for calculating preferences is detailed in Appendix F.4, Algorithm 1. Four values of $a$, $b$, $c$, and $d$ are computed to support the final steerability calculation as represented in Figure 5. For example, $a$ is as the difference between the base preference and zero. A value of zero corresponds to a model fully steered toward Safety, and one corresponds to a model fully steered toward Self-Esteem. Using these four values, steerability toward Safety is defined as $\frac{b}{a}$, while steerability toward Self-Esteem is given by $\frac{c}{d}$.
Findings.
@@
We collect base preference and difference of steerability across all value pairs and models and calculate the correlation. Our analysis reveal a significant negative correlation (Spearman’s correlation coefficient (Spearman, 1961), r=0.243,p<0.005r=-0.243,p<0.005), suggesting that strong inherent preference for one value hinders the ability to steer towards the opposing value. When considering individual model families, we observe nuanced results. Specifically, Qwen and Llama showed statistically significant negative correlations of -0.629 and -0.547, respectively, while Mistral, GPT-4o, and Claude-3.5 did not exhibit statistically significant correlations.
@@
We also investigate how steerability varies between smaller and larger models. To assess this, we perform the Wilcoxon test comparing small and large models, focusing on two aspects: (1) the absolute difference in steerability between the two values within each pair, and (2) the aggregate steerability, which is the sum of steerability toward each value in a pair. For instance, consider the value pair Safety vs. Self-Esteem; if the steerability toward Safety is 0.7 and the steerability toward Self-Esteem is 0.8, the absolute difference in steerability is 0.1, and the total steerability is 1.5. The first test, on absolute differences, reveal that smaller models (Qwen2.5-7B, Llama3.1-8B, Mistral-24B, GPT-4o-mini, Claude-3.5-Haiku) are significantly more likely to show greater steerability toward one value over the other ($p<0.005$). In contrast, the second test shows that larger models (Qwen2.5-72B, Llama3.3-70B, Mistral-123B, GPT-4o, Claude-3.5-Sonnet) exhibit significantly higher overall steerability ($p<0.0001$). Per-value steerability comparison results can be found in Appendix F.5.
4.4 Question Framing Comparisons via First- vs. Third-Person Perspectives
We conduct a comparative analysis of question framing to assess the influence of first-person versus third-person perspectives. Our default prompt adopts a third-person perspective, as we think LLMs are commonly used to support human decision-making process. However, drawing on the concept of self-distancing (Grossmann & Kross, 2014), first-person framing may make scenarios feel more personally relevant to the model, potentially leading to differences in response.
To study this, we compare performance across Straightforward, Simple Contrast, and Swayed Contrast categories, and observe a significant decline in performance when first-person perspectives are used (Wilcoxon test, $p<0.001$). We further analyze the reasons of the performance drop using steerability based on value pairs (Figure 35 in Appendix F.2). The analysis reveal that most settings exhibit increased steerability when framed from a third-person perspective, except for several pairs colored in green. Notably, first-person framing proved particularly effective when steering toward Safety relative to Self-Esteem, likely because Safety was strongly emphasized during model training. The first-person framing aligns more naturally with this inherent preference towards Safety, enhancing its effectiveness.